OpenStudy (anonymous):

Jane buys a dress for $25.98 and a blouse of $ 14.99 the sales tax is $2.54 A) what is the total cost of her purchase B) If jane buys with a $50 bill how much change should she recieve?

OpenStudy (mokeira):

50-(25.98+14.99+2.54) 50-43.51 =$6.49

OpenStudy (anonymous):

Thank you soooo much!!

OpenStudy (anonymous):

A) 25.98+14.99+2.54=43.51 total cost is 43.51 subrtract that from 50 for 6.49 which is the change she gets

OpenStudy (anonymous):

43.51

OpenStudy (anonymous):

so is the answer 6.49 or 43.51?

OpenStudy (anonymous):

both answer to A: 43.51 answer to B: 6.49

OpenStudy (anonymous):

Ohh OK so got thank you can you help me with one more?

OpenStudy (anonymous):

yea?

OpenStudy (anonymous):

OK what is 296.0/0.08?

OpenStudy (anonymous):

/= divided

OpenStudy (mokeira):

is this what you mean\[\frac{ 296 }{ 0.08 }\]

OpenStudy (anonymous):

no i mean to divied

OpenStudy (anonymous):

yea thats what @Mokeira ment to it is 3700

OpenStudy (anonymous):

OH OK thannks

OpenStudy (mokeira):

lol... hope i did not confuse you

OpenStudy (anonymous):

no you did not lol i'm only in 7th grade you know?

OpenStudy (mokeira):

dont worry...practice will make this all clear in time...dont let the class define your skills lol

OpenStudy (anonymous):

haha that level is the easy part enjoy it while it lasts

OpenStudy (anonymous):

thanks hey one more?

OpenStudy (anonymous):

go on then

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

Given: np m3 = 62° 4 and 12 are complementary The measure of 11 equals 107°. 73°. 45°. 63°.

OpenStudy (anonymous):

wait there is a drwaing that comes with it

OpenStudy (anonymous):

ok i was wondering :3

OpenStudy (anonymous):

here you go

OpenStudy (anonymous):

oh wow ok one sec

OpenStudy (anonymous):

OK thanks you know i'm in 7th grade it feelsm i'm in high schooll :O:O

OpenStudy (anonymous):

yea u said, im in 12th grade now

OpenStudy (anonymous):

the answers are 107 degress 73 dergress 45 degress 63 degress

OpenStudy (anonymous):

okay so when it says that 12 and 4 are complementary it means that the angles in those spaces are exactly the same

OpenStudy (anonymous):

OK.......

OpenStudy (anonymous):

so do i like multiply

OpenStudy (anonymous):

but thats all ik from that sorry i struggle a bit with math so maybe close this question and write this one out as new

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks form trying

OpenStudy (anonymous):

np

OpenStudy (anonymous):

t2ul